What would be the pressure of 1.5 moles of Oxygen gas in a vessel with a volume of 5.0 L and temperature of 24 DEGREES CELCIUS?


1
Expert's answer
2022-04-13T10:49:04-0400

P = ???

V = 5.0 L

n = 1.5 mol

R = 0.0821 atm L mol−1 K−1 

T = 24°C + 273.15 = 297.15 K


Solution:

The ideal gas law relates the amount of gas present to its pressure, volume, and temperature.

The ideal gas law is written as PV = nRT, 

where P is the pressure, V is the volume, n is the number of moles of gas, R is the ideal gas constant, and T is the temperature.


Rearrange the equation to solve for P:

P = nRT / V

Plug in the numbers:

P = (1.5 mol × 0.0821 atm L mol−1 K−1 × 297.15 K) / (5.0 L) = 7.32 atm

P = 7.32 atm


Answer: The pressure of oxygen gas would be 7.32 atm